OpenStudy (anonymous):

Find all odd primes which \(\Large \frac{2^{p-1}-1}{p} \) would be a perfect square number.

OpenStudy (anonymous):

not a very easy question :P

OpenStudy (anonymous):

okk so \(2^{p-1}-1 =0 \mod p\) we need to find p that satisfy \(2^{p-1}-1=p*n^2\)

OpenStudy (anonymous):

i had a feeling of there is no p that makes it true ( still my conjecture not sure :P )

OpenStudy (anonymous):

no there are some numbers ;)

OpenStudy (anonymous):

awww cute , then lemme work a bit

OpenStudy (anonymous):

aww 7 work

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and 3

OpenStudy (anonymous):

brb

ganeshie8 (ganeshie8):

Deja Vu! I'll let @Marki figure this out :)

OpenStudy (anonymous):

since when u learned french :O

OpenStudy (anonymous):

ok n^2=1 mod 8 since its odd perfect square

OpenStudy (anonymous):

\(2^{p-1}-1=p*n^2=p\mod 8\) so we would have \(2^{p-1}=p+1\mod 8\) idk if this work

OpenStudy (anonymous):

oh since its odd i should think of 1 mod 2

OpenStudy (anonymous):

hmm got it ! so only 3,7 nice

OpenStudy (anonymous):

not a nad start! ;) but prove that only 3,7 would work...

OpenStudy (anonymous):

yeah , the only primes i guess there is no more

OpenStudy (anonymous):

i'll show u some other time not feeling like doing NT :3

OpenStudy (mathmath333):

\(\)

OpenStudy (anonymous):

haha @mathmath333 style :P

OpenStudy (mathmath333):

just the bookmark

OpenStudy (anonymous):

hmm,a little help: \(\Large \frac{2^{p-1}-1}{p} = \frac{ (2^{\frac{ p-1 }{ 2 }}-1)(2^{\frac{ p-1 }{ 2 }}+1) }{ p }\)

OpenStudy (freckles):

*

OpenStudy (turingtest):

I remember back when I used to be able to do harder problems than most other user on OS. Those days are (thankfully) long gone, as threads like these remind me :) *bookmark

OpenStudy (turingtest):

users*